The length of a rectangle is six times its width. If the perimeter of the rectangle is 112 in. , find its area.
Answer by mananth(16946)

Perimeter = 112
Length 6x
Width =x cm

Perimeter = 2(6x+x)
Perimeter = 14x
112 = 14x x=
8 in Width
Length = 6*8=48 in
Area = 48*8=384 sq. in